OpenStudy (parissfrance09):

Help me so confused :( Simplify. 3^2 · 3^4 · 3^6 A) 3^0 B) 3^10 C) 3^12 D) 3^48

OpenStudy (parissfrance09):

@babyx3boo @Helpppmeee1999 @nonon @sleepyhead314

OpenStudy (anonymous):

3^12 C

OpenStudy (parissfrance09):

can you tell me how you got that answer?

OpenStudy (parissfrance09):

@hodori0719

OpenStudy (sleepyhead314):

property of exponents x^n * x^m = x^(n+m) for example 2^3 * 2^5 = 2^(3+5) = 2^8

OpenStudy (sleepyhead314):

so 3^2 · 3^4 · 3^6 = 3^(2+4+6) = 3^

OpenStudy (parissfrance09):

ohhh i undertsand now thanks